# Entanglement-enabled advantage for learning a bosonic random displacement channel

### v1

@@ Abstract
[…]
Thus, when $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, $T_{b}=1-O(1/n)$, $T_{a}=1-O(1/n)$ and $r=\Omega(\log n)$, the sample complexity becomes $N=O(\epsilon^{-2}\log\delta^{-1})$ as in the lossless case. For practically relevant squeezing and including loss prior to Bell measurement, we compare the sample complexity for the lossy TMSV+BM protocol and the lossless entanglement-free lower bound in Fig. 3, finding a significant entanglement-enabled advantage in realistic experimental settings. Specifically, for reasonable parameter choices such as squeezing parameter $r=1$, loss rate $10\%$, and $\kappa=O(1)$, we can achieve a factor of $10^{4}~{}(10^{8})$ advantage for around $n=30~{}(60)$ modes. Although the $10^{9}$ number of samples required to achieve the advantage seems large, the state-of-the-art quantum optics experiments (e.g., Refs. [36, 37]) can attain such number of samples in a reasonable time with high sampling rate up to 160 GHz.
Discussion.— We proved that schemes that exploit entanglement with an ancillary quantum memory can learn $n$-mode bosonic random displacement channels with exponentially fewer samples compared to entanglement-free schemes. Our results show that the information-theoretic framework for learning studied in DV quantum systems [17, 19] can be generalized to the CV setting and have powerful implications. We anticipate that these techniques can be applied to other CV learning tasks as well. In addition, our analysis suggests that the separation in sample complexity between entanglement-assisted and entanglement-free protocols may be realized in the near future.
Apart from their theoretical interest, random displacement channels can also be practically relevant in, e.g., modeling noise in bosonic systems. As in the qubit case [38], we expect that noise tailoring methods can transform more general noise models into random displacement channels; therefore efficiently learning random displacement channels can be useful for benchmarking CV quantum systems [39, 40] and for error mitigation.
Displacement estimation is also studied in quantum metrology (see e.g. [41, 42, 43]). A task often considered in metrology is learning an unknown unitary displacement or phase transformation acting independently on each mode [44, 42, 36, 45, 46, 47] whereas the task analyzed in this word is learning an unknown mixture of multimode displacements. Furthermore, while the goal in metrology is typically to learn one or a few parameters, in our case, the parameter space is very large. Therefore, the methodology in the two settings is quite different. Connections between metrology and bosonic channel learning are worthy of further exploration.
@@ Acknowledgements.
We thank Mankei Tsang, Yuxin Wang, Ronald de Wolf, Mingxing Yao, Ming Yuan for insightful discussions. C.O., S.C., Y.W., L.J. acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209), NSF (OMA-1936118, ERC-1941583, OMA-2137642), NTT Research, Packard Foundation (2020-71479). J.P. acknowledges support from the U.S. Department of Energy Office of Science, Office of Advanced Scientific Computing Research (DE-NA0003525, DE-SC0020290), the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the National Science Foundation (PHY-1733907). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center. S.Z. acknowledges funding provided by the Institute for Quantum Information and Matter and Perimeter Institute for Theoretical Physics, a research institute supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. J.A.H.N, Z.L., J.S.N and U.L.A acknowledge support from DNRF (bigQ, DNRF142), IFD (photoQ) and EU (CLUSTEC, ClusterQ ERC-101055224, GTGBS MC-101106833).

### v2

@@ Abstract
For the same task as in Theorem 1, a TMVS+BM scheme with squeezing parameter $r$ and transmission rates before and after the channel $T_{b}$ and $T_{a}$ respectively, can estimate any $\lambda(\beta)$ to error $\epsilon$ with success probability $1{-}\delta$ using the number of samples $N=8e^{2e^{-2r_{\text{eff}}}|\beta|^{2}}\epsilon^{-2}\log 4\delta^{-1}$, where $r_{\mathrm{eff}}$ is defined according to Eq. (14).
Thus, when $|\beta|^{2}\leq\kappa n$ with a constant $\kappa>0$, $T_{b}=1-O(1/n)$, $T_{a}=1-O(1/n)$ and $r=\Omega(\log n)$, the sample complexity becomes $N=O(\epsilon^{-2}\log\delta^{-1})$ as in the lossless case. For practically relevant squeezing and including loss prior to BM, we compare the sample complexity for the lossy TMSV+BM protocol and the lossless entanglement-free lower bound in Fig. 3, finding a significant entanglement-enabled advantage in realistic experimental settings. Specifically, for reasonable parameter choices such as squeezing parameter $r=1$, loss rate $10\%$, and $\kappa=O(1)$, we can achieve a factor of $10^{4}~{}(10^{8})$ advantage for around $n=30~{}(60)$ modes. Although the $10^{9}$ number of samples required to achieve the advantage seems large, the state-of-the-art quantum optics experiments (e.g., Refs. [46, 47]) can attain such number of samples in a reasonable time with high sampling rate up to 160 GHz.
Discussion.— We proved that schemes exploiting entanglement with an ancillary quantum memory can learn $n$-mode random displacement channels with exponentially fewer samples compared to entanglement-free schemes. Our proof technique generalizes the information-theoretic framework for learning studied in DV quantum systems [17, 19] to the CV setting. We anticipate that many other results concerning DV systems can be generalized using similar methods to CV learning tasks, such as learning CV quantum states or other CV channels. Also, our analysis suggests that the separation in sample complexity between entanglement-assisted and entanglement-free protocols may soon be realized experimentally.
Besides their theoretical interest, random displacement channels can be practically relevant in, e.g., modeling noise in bosonic systems. As in qubit cases [48], we expect that noise tailoring methods can transform more general noise models into random displacement channels; therefore efficiently learning random displacement channels can be useful for benchmarking CV quantum systems [49, 50, 39] and error mitigation.
Displacement estimation is also studied in quantum metrology (see e.g. [51, 52, 53]). A task often considered in metrology is learning an unknown unitary displacement or phase transformation acting independently on each mode [54, 52, 46, 55, 56, 57] whereas the task analyzed in this work is learning an unknown mixture of multimode displacements. Furthermore, while the goal in metrology is typically to learn one or a few parameters, in our case, the parameter space is very large. Therefore, the methodology in the two settings is quite different. Connections between metrology and bosonic channel learning are worthy of further exploration.
@@ Acknowledgements.
We thank Mankei Tsang, Yuxin Wang, Ronald de Wolf, Mingxing Yao, Ming Yuan for insightful discussions. C.O., S.C., Y.W., L.J. acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209), NSF (OMA-1936118, ERC-1941583, OMA-2137642), NTT Research, Packard Foundation (2020-71479). J.P. acknowledges support from the U.S. Department of Energy Office of Science, Office of Advanced Scientific Computing Research (DE-NA0003525, DE-SC0020290), the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the National Science Foundation (PHY-1733907). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center. S.Z. acknowledges funding provided by the Institute for Quantum Information and Matter and Perimeter Institute for Theoretical Physics, a research institute supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. J.A.H.N, Z.-H.L., J.S.N-N. and U.L.A acknowledge support from DNRF (bigQ, DNRF142), IFD (PhotoQ, 1063-00046A) and EU (CLUSTEC, ClusterQ ERC-101055224, GTGBS MC-101106833). This research was supported by Quantum Technology R&D Leading Program (Quantum Computing) (RS-2024-00431768) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT (MSIT))